OpenStudy (anonymous):

Suppose the population of a town is 100,000 in 1999. The population increases at a rate of 4.5% every year. What will be the population of the town in 2005. Round answer to the nearest whole number.

OpenStudy (lucaz):

let P be the initial population, r is the rate for every year in 1 year = P+P(r) = P(1+r) now the current population is P(1+r) in 2 years= P(1+r)+P(1+r)r = P(1+r)(1+r) = P(1+r)^2 and it goes on the same way.. from 99 to 2005 = 6 years then you have P(1+r)^t where: P=100,000 r=4.5/100=0.045 t=6